THE MICROWAVE SPECTRUM OF THE HCOOCD_{2}H SPECIES OF METHYL FORMATE^{a}

\underline{L. H. COUDERT}, LISA, CNRS/Universités Paris Est et Paris Diderot, 61 Avenue du Général de Gaulle, 94010 Créteil, France; T. R. HUET, L. MARGULÈS, R. MOTIYENKO, Laboratoire PhLAM, UMR 8523 CNRS, Bât. P5, Université des Sciences et Technologies de Lille 1, 59655 Villeneuve d’Ascq Cedex, France; and H. MOLLENDAL, Centre for Theoretical and Computational Chemistry (CTCC), University of Oslo, P. O. Box 1033, Blindern, 0315 Oslo, Norway.

Methyl formate is a non-rigid molecule displaying internal rotation of its methyl group. The microwave spectra of its normal^{b} and mono deuterated HCOOCH_{2}D^{c} species have already been studied and values for the tunneling splitting due to the internal rotation were determined. The normal species displays a 405 MHz A/E splitting, the mono deuterated one, a smaller 84.76 MHz A′/A′′ splitting. For the bideuterated species HCOOCD_{2}H, the value of this splitting is not known as its microwave spectrum has not been studied yet. In this paper experimental and theoretical investigations of the microwave spectrum of HCOOCD_{2}H are presented. More than 9000 transitions were measured with a submillimeter wave spectrometer. About 20 lines were recorded with a molecular beam spectrometer. Like for the mono deuterated species,^{c} depending on the location of the only hydrogen atom of the methyl group, two configurations arise. The C_{s}-symmetry H-in plane configuration displays a rigid rotator spectrum and its data was analyzed using a Watson-type Hamiltonian. The C_{1}-symmetry H-out of plane configuration undergoes the large amplitude internal rotation. Its data was analyzed using the so called water dimer formalism^{d} which allowed us to accurately reproduce the observed frequencies and to obtain the value of the tunneling splitting as well as the parameters involved in its rotational dependence. The hyperfine structure due to quadrupole coupling at the two deuterium atoms was also analyzed. Unexpectedly, for the H-out of plane configuration, the observed hyperfine patterns are neither those expected for two equivalent deuterium atoms nor those of a rigid molecule.

aThis work is supported by the ANR-08-BLAN-0054 and ANR-08-BLAN-0225 contracts.

bIlyushin, Kryvda, and Alekseev, J. Mol. Spec. 255 (2009) 32.

cMargulès, Coudert, Møllendal, Guillemin, Huet, and Janečkovà, J. Mol. Spec. 254 (2009) 55.

dHougen, J. Mol. Spec. 114 (1985) 395; and Coudert and Hougen, J. Mol. Spec. 130 (1988) 86.